# Preserving Multilingual Quality While Tuning Query Encoder on English Only

### v1

@@ Abstract
A dense passage retrieval system can serve as the initial stages of information retrieval, selecting the most relevant text passages for downstream tasks. In this work we conducted experiments with the goal of finding how much the quality of a multilingual retrieval could be degraded if the query part of a dual encoder is tuned on an English-only dataset (assuming scarcity of cross-lingual samples for the targeted domain or task). Specifically, starting with a high quality multilingual embedding model, we observe that an English-only tuning may not only preserve the original quality of the multilingual retrieval, but even improve it.
@@ 1 Introduction
Advances in neural NLP methods have resulted in high quality dense vector text representations Reimers and Gurevych (2019); Cer et al. (2018); Conneau et al. (2017). Such representations are often used at the initial stages of an information retrieval system, selecting the most relevant documents, ranked relative to the query Xiong et al. (2020); Zhan et al. (2020, 2021); Ren et al. (2021b). A dual encoder is successfully used to train the representations Karpukhin et al. (2020); Ren et al. (2021a); Qu et al. (2021); Hofstätter et al. (2021); Ni et al. (2022); Dong et al. (2022). A dual encoder dense passage retrieval system is efficient for two main reasons: (1) it allows using simple inner product of query and document representations, and (2) it allows modifying the query representation for a task or domain, while keeping the stored and precomputed (query-invariant) document representations intact.
If the representation was pretrained in a multilingual setting, tuning on English-only samples may be expected to degrade the multilingual qualities and there may not be enough cross-lingual samples for tuning on a specific domain or types of queries. A multilingual query generator may be employed to overcome a shortage of cross-lingual data Ren et al. (2022); Zhuang et al. (2023), but, in this work, we follow an arguably simpler strategy. In order to understand the effect of English-only tuning on multilingual qualities of a representation, and to assess a possible degradation, we consider a simple setup: A state of the art multilingual embedding model is taken as the starting point, and fine-tuned by English only samples as a query representation part of the dual encoder.
@@
    We show that fine-tuning of the query encoder on an English-only dataset may not only preserve multilingual qualities of query-document embeddings matching, but even improve them.
@@
    We consider tuning with parts of query encoder frozen, and suggest a conjectural explanation of the observations.
@@ 2 Setup
2.1 Model
In what follows, we use a state-of-the-art multilingual model intfloat/multilingual-e5-small [^1] https://huggingface.co/intfloat/multilingual-e5-small Wang et al. (2024) which we will call here $E5$. The model provides embeddings of a practical small size of $384$. For most of the evaluations, we also consider results using sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2222https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019), referred to as $L12$; this model provides embeddings of the same size $384$.
@@ 2.2 Datasets
Here we made use of the following datasets:
  1. 1.
    MSMARCO Triplets [^3] https://huggingface.co/datasets/sentence-transformers/embedding-training-data/blob/main/msmarco-triplets.jsonl.gz, which we split on training, validation and test parts.
  2. 2.
    ARXIV subset, which we take from a mirror of arxiv (version 173) [^4] https://huggingface.co/datasets/arxiv-community/arxiv_dataset [^5] https://www.kaggle.com/datasets/Cornell-University/arxiv
  3. 3.
    XNLI multilingual dataset [^6] https://huggingface.co/datasets/facebook/xnli Conneau et al. (2018), of which we use the test subset.
The MSMARCO dataset consists of 499184 samples, with each sample being a tuple given as (query, positives, negatives). The “positives” are the correct answers to the query, and the “negatives” are semantically similar, but incorrect answers. For most samples there is only one positive, but many negatives. For our tuning we simply select the very first positive and the very first negative. Thus, each sample gives one triplet (anchor, positive, negative) for contrastive learning, where the query is taken as an anchor. We keep the first 487983 samples (or 34856 batches if each batch is 14 triplets) for tuning, leaving the next 4200 samples (300 batches) for validation, and the last 7000 samples for evaluation. During evaluation we create all possible triplets from the 7000 samples (using all positives and negatives); this makes 357642 evaluation triplets.
We made two flavors of evaluation arxiv triplets from this arxiv subset. In the first version, the anchor is the title, the positive is the corresponding abstract, and the negative is another random abstract. In the second version the anchor is the first sentence of an abstract, the positive is the rest of the abstract, and the negative is a similar piece (first sentence excluded) of another, random (with fixed seed) abstract. (For full reproducibility of obtaining arxiv triplets, see Appendix A.)
In evaluating on triplets (of MSMARCO and ARXIV) we compare the similarity (or distance) between the anchor and the positive vs the negative. If the positive does not turn out to be closer than the negative to the anchor, we count this as an error.
XNLI consists of pairs of sentences which are human-labeled as entailment, neutral or contradiction. The test subset (which we use) contains 1670 pairs for each of these labels and each sentence is presented in 15 languages: [’ar’, ’bg’, ’de’, ’el’, ’en’, ’es’, ’fr’, ’hi’, ’ru’, ’sw’, ’th’, ’tr’, ’ur’, ’vi’, ’zh’]. We use 225 versions of the pairs, because each sentence of the pair can be in any of the 15 languages. At evaluation the first sentence serves as the query (the embedding is taken by the query model), and the second one as the text. We expect that the sentences of an entailment pair should be closer to each other than the sentences of any neutral pair, or of any contradiction pair. Whenever this does not happen, we count this as an error.
data Evaluation $E5$ $L12$
MM N tot 357642
N errors (cos) 16688 54156
N errors (dist) 17036 55103
ARX-F N tot 253140
N errors (cos) 4014 12488
N errors (dist) 3939 17009
ARX-T N tot 253140
N errors (cos) 399 3547
N errors (dist) 453 4264
XNLI N total 2788900
entail-neutral (cos) 10.8% 10.2%
entail-contr (cos) 10.0% 7.2%
entail-neutral (dist) 10.5% 10.1%
entail-contr (dist) 9.6% 7.8%
Table 1: The count of errors for the original untuned models $E5$ and $L12$, on the datasets noted in the first column: MM - MSMARCO last 7000 samples (357642 triplets, see Section 2.2); ARX-F - arxiv subset with the abstract’s first sentence as an anchor; ARX-T - arxiv subset with the title as an anchor; XNLI - XNLI test subset providing 1670x1670=2788900 comparisons of entailment pairs vs neutral pairs (and the same amount of entailment pairs vs contradiction pairs). For XNLI the errors are averaged over 225 (15x15) language-language versions, and shown as percent of $Ntotal$.
2.3 Tuning
We tune a dual encoder by contrastive learning on MSMARCO in the following simple regime:
  1. 1.
    The text encoder is fully frozen; some parts of the query encoder can be frozen.
  2. 2.
    Unless specified otherwise, the batch size is 14, the learning rate is 5e-8 and the contrastive learning margin is $0.1$. The loss is defined by the triple margin loss.
  3. 3.
    Unless specified otherwise, there are $1000$ batches per epoch, and stopping occurs after $10$ consecutive non-improvement epochs. The improvement is measured on the validation subset after each epoch. The model is considered to be improved if (on the validation subset) both the loss and the count of errors are decreased.
We have extensively considered changing to this tuning parameters in Appendix  B (and in Appendix  C for model $L12$); variations of tuning scheduling are considered in Appendix  D.
@@ 3 Observations
3.1 Original multilingual embedding
To establish a baseline before any fine-tuning, and to ensure our evaluation is not too easy, we measure the errors of the original $E5$ model on the data described in Section 2.2 and show the results in Table 1. (We also measure the errors of $L12$, which does far worse on English data (MSMARCO, ARXIV) by comparison.)
The count of errors on the triplets (MSMARCO, ARXIV) is straightforward: it is an error when a positive is not closer than a negative to the anchor of the triplet. On XNLI we sum up the error count over all language-language pairs and divide the sum by the number (255=15x15) of such pairs. This averaged error is shown as a percentage of the total (2788900) comparisons; each comparison here is either a comparison of an entailment-labeled sample with a neutral-labeled sample (entail-neutral in the table) or a comparison of an entailment-labeled sample with a contradiction-labeled sample (entail-contr in the table). An error was counted whenever the sentences of an entailment sample happened to be farther from each other than the sentences of a neutral (or contradiction) sample.
The amount of errors (Table 1) is reasonable enough to consider how tuning would affect them.
@@ 3.1 Original multilingual embedding
- 7.47 8.46 4.51 4.85 1.75 5.52 222/0 215/2 194/4 147/23
emb.base 7.32 8.82 4.16 5.08 3.51 7.73 222/0 217/1 201/2 159/21
emb 7.30 8.82 4.21 5.05 3.51 7.73 222/0 217/1 201/2 158/20
emb, B0a 7.30 8.76 4.09 5.00 3.26 7.73 222/0 217/1 200/2 159/21
emb, B0a,i 7.48 9.00 4.36 5.03 3.26 7.73 223/0 219/0 199/2 156/25
emb, B0a,i,od 7.31 8.82 4.09 4.85 3.51 7.73 222/0 217/1 200/2 158/21
emb, B0 7.35 8.78 4.09 5.10 3.51 7.73 222/0 217/1 200/2 159/19
emb, B0-5 7.87 9.39 5.11 5.74 3.26 7.51 219/0 213/3 200/5 157/25
emb, B0-10 1.45 2.57 0.17 0.86 0.00 0.44 123/0 112/0 21/0 25/10
Table 2: Evaluations of the $E5$ query model tuned on MSMARCO as described in Section 2.3. The rows are in the order of increased freezing (at tuning): from no freezing (top row) to freezing everything up to the last transformer block (the last row); the last block is $B11$. The $emb.base$ model has only the first three layers of the embedding block frozen (tokens, positions, token-types). The $emb$ model has the full embedding block frozen. The other notations: $B0$ is the full first transformer block; $B0$-$5$ are the first 6 blocks; the extensions $a$, $i$, $od$ (for $B0$) correspond to the layers $attention$, $intermediate$ and $output.dense$ of the block. The columns $c\%$ and $d\%$ show the improvement (in percents) relative to the original $E5$ model by evaluation with cosine ($c$) and distance ($d$). The columns $c$+/- and $d$+/- show how many language pairs are improved ($+$) or worsened (--) by statistically significant change of the number of errors, measured with cosine ($c$) or distance ($d$). Values that did not pass two-tailed test are gray-colored.
@@ 3.2 Tuning partially frozen query model
Tuning (as specified in Section 2.3) of the dual encoder with the text encoder frozen and query model free ends up with the result shown in the first row of Table 2. The improvement $I$ of the performance on triplets (columns $c\%$ and $d\%$ in the table) is presented by the ratio
$I=\frac{N_{e}-\tilde{N_{e}}}{N_{e}}$ (1)
where $N_{e}$ is the original number of errors, and $\tilde{N_{e}}$ is the number of errors after the tuning. The improvement $I$ is shown as a percentage and is measured either by cosine (columns $c\%$) or by distance (columns $d\%$).
The improvement on XNLI is shown by the number of language pairs in which the number of errors changed in a statistically significant way. In the columns c+/- (cosine similarity is used) and d+/- (distance is used) the number of improved language pairs is shown in the “numerator” (+), and the number of worsen pairs is shown in the “denominator” (-) in the cell. The change for a language-language pair is considered statistically significant if the Z-value for two-tailed test exceeds $1.96$ (which corresponds to 0.025 on each tail).
For example, the first row of Table 2 shows that the tuning of the query model with no frozen parts improves its cosine-measured performance on the MSMARCO evaluation samples by $7.47\%$; it also improves on ARXIV, and, interestingly, it has shown more improvement than worsening on multilingual XNLI data. For cosine-assessed comparisons of entailment-labeled and neutral-labeled XNLI samples, the tuning statistically significantly improved 222 out of 225 language pairs, without worsening a single one. Of the XNLI columns the worst result is in the last one: the distance-assessed comparisons of entailment-labeled and contradiction-labeled pairs. It still shows more improved than worsened language pairs (147 vs 23).
In the next rows of Table 2 we increase the frozen part of the query model, guided by the intuition that the lowest, closest to input, layers (especially the embedding block) may be dealing with specifics of the languages, while the higher transformer blocks may be less dependent on the languages. Indeed, as we freeze the lower three layers of the embedding block (word, position and token type embedding layers), we see even better improvement on XNLI. The third row $emb$ corresponds to freezing the whole embedding block i.e. freezing its $LayerNorm$ in addition to $emb.base$); this gives similar results.
Freezing beyond the embedding block in the query model, we freeze the embedding block together with parts of the first transformer block, such as its attention layers ($B0_{a}$), then also its intermediate layers ($B0_{a,i}$), then also output dense layers ($B0_{a,i,od}$), and finally the whole block $B0$. The XNLI improvement remains about as good as when only the embedding block was frozen ($emb$ row). Freezing half of the transformer ($emb$ and $B0$-$5$) still gives similar results, if only slightly worse on XNLI. However, freezing everything up to the last transformer block (last row in Table  2) naturally diminishes the overall effect of the tuning. For verifying our intuition, we also tried the opposite to freezing the embedding: we had frozen everything except the embedding block. In such setting our tuning regime could not improve the model at all.
In Appendix E we show improvements for each XNLI language pair, for the case of our default tuning (Section 2.3), with embedding block frozen.
Our results do not depend strongly on the tuning parameters (Appendix B), unless the parameters are changed too much, especially the contrastive learning triple loss margin (Appendix B.3). Naturally, increasing learning rate delivers more gains on MSMARCO and ARXIV, while slightly reducing gains on XNLI. The learning rates between 2e-8 and 6e-8 may be considered as the best balance, but a rate as high as 1e-7 may be preferred, depending on languages of interest, see Appendix B.1.
The batch sizes as different as 7, 14, 28, 56 and 112 deliver similar gains, if the number of samples per epoch 14000 is kept the same (Appendix B.2, Table 4). The batch size 112 is ’safer’, giving less gains but not degrading any language pairs if evaluated with cosine similarity, and only single one when evaluated with distance. If the number of batches 1000 (rather than number of samples) is kept the same (Table 5 then the large batch size 112 delivers highest gain on MSMARCO, but far worse results on ARXIV and XNLI.
The margin of contrastive learning is good between 0.05 and 0.15 (Appendix B.3). Increasing it beyond 0.15 diminishes gains in all the datasets, and at margin 0.5 the number of degraded language pairs becomes comparable to the number of improved language pairs.
Varying our stopping criterion (10 consecutive non-improvement epochs) between 5 and 15 does not have much effect (Appendix B.4). Interestingly, switching to a less straightforward tuning regime, i.e. using scheduling and weight decay may deteriorate the performance gains on MSMARCO and ARXIV, and may degrade performance on many language pairs in XNLI.
The model $L12$, having much more errors on English samples at start (Table 1), also can improve/preserve performance on XNLI, but not as good as $E5$ (Appendix C). We speculate that this may be related to its worse starting point.

### v3

@@ Abstract
A query encoder of a dual passage retrieval system can be tuned for specific types of queries or domains, while the precomputed and stored documents representations are kept intact. Switching from one query encoder to another when needed is easily feasible, unlike overhauling the embeddings of a whole knowledge base. In this work we raise a question: Can the generic, original qualities of the encoder be preserved or at least left not too degraded when it is tuned on a narrow domain? We conducted experiments on a high quality multilingual embedding model: Tuning it on a single English-only dataset, we observe that the tuning not only preserves the multilingual qualities, but even improves them. The embedding qualities on distinctly different data are also improved or at least preserved. Drawing on our observations, we suggest a more general hypothesis: Tuning with intentionally low learning rate can preserve or improve a system’s properties acquired in training, but not specifically targeted by tuning. We call this adiabatic tuning and provide tentative explanations.
@@ 1 Introduction
Advances in neural NLP methods have resulted in high quality dense vector text representations Reimers and Gurevych (2019); Cer et al. (2018); Conneau et al. (2017). Such representations are often used at the initial stages of an information retrieval system, selecting the most relevant documents, ranked relative to the query Xiong et al. (2020); Zhan et al. (2020, 2021); Ren et al. (2021b). A dual encoder is successfully used to train the representations Karpukhin et al. (2020); Ren et al. (2021a); Qu et al. (2021); Hofstätter et al. (2021); Ni et al. (2022); Dong et al. (2022). A dual encoder dense passage retrieval system is efficient for two main reasons: (1) it allows using the simple inner product of query and document representations, and (2) it allows modifying the query representation for a task or domain, while keeping the stored and precomputed (query-invariant) document representations intact.
If the representation was pretrained in a multilingual setting, tuning on English-only samples may be expected to degrade the multilingual qualities and there may not be enough cross-lingual samples for tuning on a specific domain or types of queries. A multilingual query generator may be employed to overcome a shortage of cross-lingual data Ren et al. (2022); Zhuang et al. (2023), but, in this work, we follow an arguably simpler strategy. In order to understand the effect of English-only tuning on multilingual qualities of a representation, and to assess a possible degradation, we consider a simple setup: A state of the art multilingual embedding model is taken as the starting point, and fine-tuned by English only samples as the query representation part of a dual encoder.
We assume that our observations of the degradation or preservation of the multilingual qualities may be generalized to other pretrained system qualities that are not directly targeted in tuning. In order to obtain preliminary confirmation of this hypothesis, we also observe the effect of tuning on the embedding quality for queries and text chunks of very different styles, the likes of which could be present in the training of the original encoder, but certainly not targeted in tuning.
@@
    We show that fine-tuning a query encoder on an English-only dataset may not only preserve multilingual qualities of query-document embeddings matching, but even improve them.
@@
    We hypothesize that a tuning regime with intentionally low learning rate (far below of what is necessary to avoid overfitting) preserves or improves the properties acquired in the training, but not targeted by tuning. We call this adiabatic tuning and suggest supporting observations and conjectural explanations.
  3. 3.
    We add a dataset with graded difficulty, based on ARXIV titles and abstracts.
Although high-resource languages can be used for cross-lingual transfer Lin et al. (2019), our setting does not have such a goal: the tuning is set to improve the query part of a dual encoder on a certain dataset, with no driving mechanism for preserving or improving the other qualities of the system.
Our starting point is one of the best (for its lean size) multilingual embedding models which differs from starting with a multilingual language model and then aligning the generated embeddings for different languages Wang et al. (2022).
@@ 2 Setup
2.1 Models
In what follows, we use a state-of-the-art multilingual model intfloat/multilingual-e5-small [^1] https://huggingface.co/intfloat/multilingual-e5-small Wang et al. (2024b) which will be referred to here as $E5$. For most of the evaluations, we also consider results using sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2222https://huggingface.co/sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 Reimers and Gurevych (2019), referred to as $L12$. Finally, we confirm some observations with monolingual intfloat/e5-small-v2333https://huggingface.co/intfloat/e5-small-v2 Wang et al. (2024a), referred to as $E5e$. All these models provide embeddings of a practical small size of $384$.
@@ 2.2 Datasets
We use MSMARCO Nguyen et al. (2018) Triplets [^4] https://huggingface.co/datasets/sentence-transformers/embedding-training-data/blob/main/msmarco-triplets.jsonl.gz for tuning and evaluation. For evaluating the qualities not targeted by tuning, we use the ARXIV dataset with negatives [^5] https://huggingface.co/datasets/primer-ai/arxiv-negatives, which we made from arxiv (version 173) [^6] https://www.kaggle.com/datasets/Cornell-University/arxiv, and the test subset of the XNLI multilingual dataset [^7] https://huggingface.co/datasets/facebook/xnli Conneau et al. (2018). We also use HOTPOTQA [^8] https://hotpotqa.github.io/ Yang et al. (2018) and SQUAD [^9] https://huggingface.co/datasets/rajpurkar/squad_v2 Rajpurkar et al. (2018, 2016) for confirming some observations (Appendices CD).
Our test subset of MSMARCO contains 357642 evaluation triplets, made of 7000 samples - all the positives and negatives are used (Appendix A).
Of ARXIV we use titles and abstracts. We made two flavors of evaluation arxiv triplets: (1) arxiv-title where a title plays role of the query (anchor), and the corresponding abstract is a positive passage, and (2) arxiv-first where the first sentence of abstract is used as the query, and the rest of it is used as a positive (Appendix B). We also use narrow versions of arxiv-first in Appendix K.
2.3 Tuning and evaluations
Unless otherwise specified, we freeze the text encoder and proceed to fine-tune only the query encoder (fully or partially unfrozen) by contrastive learning on MSMARCO (or narrow ARXIV subsets, Appendix K) with a learning rate of 5e-8, batch size of 14 and the triple margin loss with margin $0.1$. Other details are in Appendix E. In our experiments we considered different settings of freezing, batch size, learning rate, the margin of triplet loss, the stopping criterion, weight decay, scheduling versions and optimizers.
In most of our evaluations, we compare the similarity (or distance) between the anchor (query) and the positive vs the negative. If the positive does not turn out to be closer than the negative to the anchor, we count this as an error. We thus characterize performance of the encoder on a query by the number of errors divided by the total number of positive-negative pairs. We call this positive-negative discrepancy (PND). The measure is easy to interpret, and its range (from 0 to 1) is the same and equally fair for any amounts of positives and negatives, as long as they exist in a selection for a query. On multiple queries we take an averaged PND. We confirm some results also using mean reciprocal rank (MRR), mean average precision (MAP) and precision at top 1 (P@1). The improvement of performance is measured as relative change of a measure $M$ (PND or MRR or other):
$I=s\frac{\tilde{M}-M}{M}$ (1)
where $M$ is for the original encoder, and $\tilde{M}$ is for the encoder after the tuning. The sign $s$=-1 for PND, because it decreases when improved, and $s$=1 for the other measures.
For evaluating XNLI we use its pairs of sentences, each sentence is given in 15 languages (Appendix F). One sentence is used as a query, another as a passage. All pairs are human-labeled as entailment, neutral or contradiction. Hence, the sentences of an entailment pair should be closer to each other than the sentences of any neutral or contradiction pair. Whenever this does not happen, we count this as an error for PND. In Appendix G we made sure that the amount of errors the original encoder makes on our datasets is large enough to consider how tuning would affect them.
@@ 3 Observations
- 7.47 8.46 5.19 5.19 1.75 5.52 222/0 215/2 194/4 147/23
emb.base 7.32 8.82 4.85 5.41 3.51 7.73 222/0 217/1 201/2 159/21
emb 7.30 8.82 4.90 5.39 3.51 7.73 222/0 217/1 201/2 158/20
emb, B0a 7.30 8.76 4.77 5.34 3.26 7.73 222/0 217/1 200/2 159/21
emb, B0a,i 7.48 9.00 5.05 5.36 3.26 7.73 223/0 219/0 199/2 156/25
emb, B0a,i,od 7.31 8.82 4.77 5.19 3.51 7.73 222/0 217/1 200/2 158/21
emb, B0 7.35 8.78 4.77 5.44 3.51 7.73 222/0 217/1 200/2 159/19
emb, B0-5 7.87 9.39 5.79 6.07 3.26 7.51 219/0 213/3 200/5 157/25
emb, B0-10 1.45 2.57 0.89 1.21 0.00 0.44 123/0 112/0 21/0 25/10
Table 1: Evaluations of the $E5$ query model tuned on MSMARCO as described in Section 2.3. The rows are in the order of increased freezing (at tuning): from no freezing (top row) to freezing everything up to the last transformer block $B11$. The $emb.base$ model has only the first three layers of the embedding block frozen (tokens, positions, token-types). The $emb$ model has the full embedding block frozen. For the other notation: $B0$ is the full first transformer block; $B0$-$5$ are the first 6 blocks; the extensions $a$, $i$, $od$ (for $B0$) denote the layers $attention$, $intermediate$ and $output.dense$ of the block. The columns $c\%$ and $d\%$ show the PND improvement (in percents) relative to the original model, accessed by cosine ($c$) or distance ($d$), grayed if not significant (Appendix I). The columns $c$+/- and $d$+/- show count of language pairs with PND significantly improved ($+$) or worsened (--).
@@ 3.1 Tuning partially frozen query model
In Table 1 we show results of tuning the dual encoder, with the text encoder frozen and query model free or partially frozen. Here and throughout the paper we use the easiest version of ARXIV (see Appendix H on performance at other levels). Freezing the embedding block appears to be the best option for preserving the multilingual qualities, and henceforth it is used unless specified otherwise. In Table 2 we confirm the improvement on six other datasets (Appendices ACD), and show some other measures.
PND MRR MAP P@1
Dataset cdcdcdcd%
MSMARCO 65 negatives 2.41 3.78 0.48 0.55 1.03 1.15 1.92 2.02
SQUAD 1.02 1.12 0.17 0.2 0.17 0.19 0.31 0.33
SQUAD min 5 0.85 1.13 0.16 0.24 0.18 0.26 0.32 0.44
HotpotQA easy 2.52 3.47 0.25 0.34 0.09 0.08 0.16 0.12
HotpotQA medium 2.53 3.57 0.33 0.49 0.07 0.09 0.11 0.13
HotpotQA hard 2.43 3.70 0.30 0.50 0.07 0.11 0.12 0.15
Table 2: Improvements for $E5$ tuned with frozen embedding block and learning rate 5e-8.
The multilingual qualities are not only preserved, but even mostly improved, especially on cosine similarity. The PND improvement is shown for each language pair separately in Figure 1. The results for the $L12$ model are similar (Appendix J). In Appendix K we also also confirm our observations with $E5$ tuned on specific categories of ARXIV.
Figure 1: Improvement of $E5$ on XNLI assessed by cosine. Query is on axis $Y$; text is on $X$.
3.2 Learning rate and adiabatic tuning
Figure 2: Evaluations on (a) XNLI and (b) the English-only datasets (MSMARCO and ARXIV) of the E5 query encoder tuned with a frozen embedding block, batch size 14, margin 0.1 using different learning rates. Values that did not pass the two-tailed test are shown with open markers.
Increasing the tuning learning rate delivers more gains on MSMARCO, while eventually reducing gains on XNLI and even ARXIV. Improvement of PND on MSMARCO and ARXIV is shown in Figure 2(b); the number of language pairs improved and degraded is in Figure 2(a). Appendix L contains the corresponding plots (Figure 10) for the fully tuned $E5$ dual encoder, and for the $L12$ and $E5e$ models. It is interesting that the $E5e$ model, not even being multilingual, still improves more than it degrades its rudimentary multilingual qualities. The effects of other tuning parameters are described in Appendix M. For example, the square-root batch size scaling rule works better than linear.
If we consider XNLI and ARXIV as indicators of how well a model keeps the learned skills while improving on narrow goals (e.g. MSMARCO), then our observation suggests there may be a slow tuning regime, at which the model preserves or even improves the existing skills which are at least a little related to the new goal. We call this adiabatic tuning, in analogy to the slow process in quantum mechanics (a system starting in an eigenstate is kept in the same evolving eigenstate). For $E5$ the learning rates between 2e-8 and 6e-8 may be considered as the best.
Our tentative explanation of adiabatic tuning is as follows: At low learning rates of tuning, the system (the encoder weights) remains in the ’minimum’ region found at pretraining. This ’minimum’ region is probably a wide well with uneven ground; the pretraining happened to terminate at some point inside the well. During tuning, the pretraining weight-space of twin encoder becomes just another surface in a family of surfaces, because of the added dimensions (the difference between the weights of the two encoders). We assume that due to continuity, the ’minimum’ region, even if being reshaped, remains a well as the query encoder weights drift away from the weights of the text encoder. Within this well, improvements of all qualities related to the former, pretraining loss, may be still correlated. But if, at high learning rate, the model is strongly modified at some iteration (i.e. by backpropagation on a particular batch), then it may move away from the well. In Appendix M.3 we present a simple attempt to identify and extend the range of adiabatic tuning.